The perimeter of a rectangle is 254 inches. The length exceeds the width by 63 inches. Find the length and the width.

Answer by mananth(16949)

let width = x
length = x+63
..
Perimeter = 2*(L+W)
2*(x+x+63)=254
2(2x+63)=254
4x+126=254
4x=254-126
4x=128
x=128/4
x=32 width
32+63=95 length
...
C H E C K
2*32+2*95=254